OpenStudy (anonymous):

how would i find this cosx= root 3 i need the value of x?

OpenStudy (amistre64):

then you need to know your basic trig angles; which is more memorization than anything else

OpenStudy (anonymous):

its not one of the basic ones

OpenStudy (anonymous):

theres a way to figure it out, but i can't remember what it is

OpenStudy (amistre64):

if you believe it isnt basic; then you need to use a calculator or a trig table; the answer is x = cos^-1(x)

OpenStudy (amistre64):

arccos is another name for the inverse function

OpenStudy (anonymous):

thats what i thought, so i have to do inverse trig to find it?

OpenStudy (anonymous):

it would be a miracle if \[\cos(x)=\sqrt{3}\]

OpenStudy (amistre64):

yes; you have to "undo" the trig :)

OpenStudy (anonymous):

when i put that in my calc it says error

OpenStudy (anonymous):

because \[-1\leq \cos(x) \leq 1\] for all x

OpenStudy (amistre64):

then the answer is undefined ;)

OpenStudy (anonymous):

yes it is not possible. the biggest cosine can be is 1 and the smallest it can be is -1

OpenStudy (anonymous):

so in particular it cannot be \[\sqrt{3}\]

OpenStudy (anonymous):

this problem has to have an answer, it was f(x)=cosx [-pi/3 , pi/3]

OpenStudy (anonymous):

don't try to find the inverse sine or cosine of anything bigger than 1 or less than -1

OpenStudy (amistre64):

unless you left out something in the problem, the answer is undefined

OpenStudy (anonymous):

that's the whole problem, but the integral evaluation of that can't be undefined

OpenStudy (amistre64):

Math is full of problems that have no answer :)

OpenStudy (anonymous):

Mevertheless, there are complex numbers whose cosines are sqrt(3)

OpenStudy (anonymous):

if this is a complex variables class ineedhelp is taking i will eat my hat

OpenStudy (anonymous):

From Mathematica 8\[x= \text{ArcCos}\left[\sqrt{3}\right]=0.+1.14622 i \]